A typical cell membrane is 8.0 nm thick and has an electrical resistivity of 1.3 ✕ 107 ·m.

If the potential difference between the inner and outer surfaces of a cell membrane is 65 mV, how much current flows through a square area of membrane 1.0µm on a side?

_____A
Answer #1

The resistance of the membrane is going to be proportional to the thickness and inversely proportional to the cross-sectional area: thus
R = 1.3e7 * 8 (nm) / (1 µm)^2
    = 1.3e7*8e-9 / (e-6)^2
    = 1.3*8 * e10
    = 1.04e11

Current I = V/R
                = 65 (mV)/(1.04e11)
                = 6.5e-2/(1.04e11)
               = (6.5/1.04)* e-13
                = 6.25e-13 A